Godgift is x years old now.
a)how old was he 7 years ago?
b)how old will he be in 15 years time from now?
c)if his age in 15 years time will be three times his age 7 years ago,how old is he now?

Answer by ankor@dixie-net.com(22740)

Godgift is x years old now.
a)How old was he 7 years ago?
x - 7
b)How old will he be in 15 years time from now?
x + 15
c)If his age in 15 years time will be three times his age 7 years ago,
how old is he now?
x + 15 = 3(x-7)
x + 15 = 3x - 21
15 + 21 = 3x - x
36 = 2x
x = 36/2
x = 18 yrs is G's present age